# Spin transport and lack of quantisation in the $A\mathrm{II}$ class on the honeycomb structure

### v1

@@ 4.2. Phase diagram of the extended Kane–Mele model
First of all, we provide the dimerisation $[H_{\mathrm{KM}}]_{{\mathcal{D}}_{i}}$. With abuse of notation, we let $T_{v}$ denote the translation operator on $\ell^{2}(\Gamma)$, whose definition is as in (2.7) with $\Gamma$ in place of $\mathcal{C}$. By straightforward computations, we obtain:
@@
[…]
@@
and when $w=-\sigma\varepsilon\lambda_{\mathrm{SO}}$. Note that on the Brillouin torus $\mathbb{T}^{2}_{*}$, we have $-k_{F}^{+}=k_{F}^{-}$.
@@
[…]
@@
with $[\,\cdot\,]_{+}:=\max\{\,\cdot\,,0\}$. Finally, defining the critical energy
@@
Note that the spin is not conserved and thus is not a good label for the energy bands. To compute the local internal gap $\Delta(k_{F}^{+})$, it is actually easier to first determine which bands are the external ones. Without loss of generality, we assume $\lambda_{\mathrm{SO}}>0$: otherwise, we could flip the sign of $w$ and $\lambda_{\mathrm{R}}$ to flip the sign of $H_{\mathrm{KM}}(k)$, the latter operation leaving the set of internal/external bands invariant. Note that $\mathcal{E}_{1,+}(k_{F}^{+})$ is always the positive external band. When $\mathcal{E}_{1,-}(k_{F}^{+})\leq\mathcal{E}_{2,-}(k_{F}^{+})$, the negative external band is $\mathcal{E}_{2,-}(k_{F}^{+})$, so that the local internal spectral gap is
@@
The condition $\mathcal{E}_{1,-}(k_{F}^{+})\leq\mathcal{E}_{2,-}(k_{F}^{+})$ is equivalent to $|\lambda_{\mathrm{R}}|\geq 2\sqrt{\lambda_{\mathrm{SO}}^{2}+|w|\lambda_{\mathrm{SO}}}$. In contrast, if $\mathcal{E}_{1,-}(k_{F}^{+})<\mathcal{E}_{2,-}(k_{F}^{+})$, $\mathcal{E}_{2,\pm}(k_{F}^{+})$ become the internal bands, and therefore $\Delta(k_{F}^{+})=2|w|$. The expression for $w_{c}$ and $\mu_{c}$ is a straightforward computation from $\Delta(k_{F}^{+})$, noticing that $\mathcal{E}_{2,+}(k_{F}^{+})$ is always an internal band. ∎
@@
We have thus shown that for $|\lambda_{\mathrm{R}}|$ small, the plot in Fig. 3 describes the quantum phase diagram of the model, consisting of three insulating phases separated by a semi-metallic one along the critical curves $w_{c}^{\pm}$. This phase diagram was already obtained for the Kane–Mele model ($r=0$) in the seminal work [KM1, Fig. 1, inset]. In particular, the central phase is a topological insulator with spin Chern number equal to $1$, whereas the outer regions have spin Chern number equal to 00. This can be directly inferred by continuity of the spin Chern number for insulators with respect to $\lambda_{\mathrm{R}}$, and thus by considering the case $\lambda_{\mathrm{R}}=0$, see (4.8), for which the spin Chern number is equal to the Chern number of $H_{+}$ in (4.8), which was computed in [Hal, Fig. 2 at $\phi=\pi/2$].
4.3. Lack of quantisation in the $\mathrm{AII}^{\hexagon}$ class
Unlike the case of charge transport, the spin conductivity as introduced in Definition 3.12 is not universally quantised for single-particle insulators.
@@ 4.3. Lack of quantisation in the $\mathrm{AII}^{\hexagon}$ class
There exist insulators $(H,\mu)$ with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ almost conserving the spin (in the sense of Definition 3.20) such that $\sigma^{\mathrm{s}}_{12}(H,\mu)\notin\frac{1}{2\pi}\mathbb{Z}$. In particular, for such insulators it holds true that
$c{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\leq\Big{|}\sigma^{\mathrm{s}}_{12}-\frac{1}{2\pi}\mbox{$S$-Chern}(P^{\mathrm{sc}})_{12}\Big{|}\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\;,$ (4.18)
@@
provided that the limit exists. Note that by Proposition 3.15, it suffices to consider only $\sigma^{s}_{12}$, as $\sigma^{s}$ is an antisymmetric tensor. Note that we could equivalently study the jump across the lower critical line $w_{c}^{-}$, compare with (4.14).
@@
Assume the contrary, that is, $\sigma^{\mathrm{s}}_{12}(H,\mu)\in\frac{1}{2\pi}\mathbb{Z}$ for any insulator $(H,\mu)$ with $H\in\mathrm{AII}^{\hexagon}$. Then, $\delta\sigma_{12}^{\mathrm{s}}(t,\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}},r)\in\frac{1}{2\pi}\mathbb{Z}$ for any value of $t,\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}},r$, which is in contradiction with (4.22).
We are left with proving the lower bound in (4.18), the upper bound being stated in Theorem 3.22. Once more, we do this by considering the pair $(H_{\mathrm{KM}},\mu_{c})$ at fixed $t,\lambda_{\mathrm{SO}},r$ non-zero, at $\lambda_{\mathrm{R}}>0$ sufficiently small and at $w$ sufficiently close to $w_{c}^{+}(\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}})$, that is, for $m\neq 0$, $m$ sufficiently small. Since all parameters but $m$ and $\lambda_{\mathrm{R}}$ are fixed, let us abridge (4.20) to $\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})$, and likewise $P^{\mathrm{sc}}=P^{\mathrm{sc}}(m)$, $w^{+}_{c}=w^{+}_{c}(\lambda_{\mathrm{R}})$, and $\delta\sigma_{12}^{s}=\delta\sigma_{12}^{s}(\lambda_{\mathrm{R}})$ (notice that $H^{\mathrm{sc}}=H_{\mathrm{KM}}|_{\lambda_{\mathrm{R}}=0}$).
@@
for $m\neq 0$ and $\lambda_{R}>0$, both sufficiently small. We observe that ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|H^{\mathrm{snc}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}=O(\lambda_{\mathrm{R}})$ (see Remark 4.2 and (3.28)), so that the claim we need to prove is that there exist an $m\neq 0$, a constant $C>0$ and an open interval $J\ni 0$, such that $|f(m,\lambda_{\mathrm{R}})|\geq C\lambda_{\mathrm{R}}^{2}$ holds true for all $\lambda_{\mathrm{R}}\in J$.
We now assume the contrary of the claim and show that this implies a contradiction. We first of all note that the contrary of the claim implies that $f(m,\lambda_{\mathrm{R}})=o(\lambda_{\mathrm{R}}^{2})$. In fact, the contrary of the claim is that for all $m\neq 0$, constants $C>0$, and open intervals $J\ni 0$, there exists at least one point $\lambda_{0}\in J\setminus\{0\}$, such that $|f(m,\lambda_{0})|<C\lambda_{0}^{2}$. Since this holds for any open interval $J\ni 0$, we can iterate the reasoning and find a sequence $(\lambda_{n})_{n}\in\mathbb{N}$ such that $\lambda_{n}\to 0$ and $|f(m,\lambda_{n})|<C\lambda_{n}^{2}$. By Remark 4.17 the limit $\lim_{\lambda_{\mathrm{R}}\to 0}f(m,\lambda_{\mathrm{R}})$ exists for any $m\neq 0$, hence $\lim_{\lambda_{\mathrm{R}}\to 0}f(m,\lambda_{\mathrm{R}})/\lambda_{\mathrm{R}}^{2}=0$ (as the constant $C$ can be chosen arbitrarily small), that is, $f(m,\lambda_{\mathrm{R}})=o(\lambda_{\mathrm{R}}^{2})$.
We define
@@
[…]
@@
$\begin{split}g(m,\lambda_{\mathrm{R}})&=\frac{1}{2\pi}+\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})\\ &=\frac{1}{2\pi}+\delta\sigma_{12}^{s}(\lambda_{\mathrm{R}})+\Big{(}\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{+},\lambda_{\mathrm{R}})\Big{)}-\Big{(}\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{-},\lambda_{\mathrm{R}})\Big{)}\\ &=-\frac{1}{2\pi}\frac{\lambda_{\mathrm{R}}^{2}r}{2t\lambda_{\mathrm{SO}}-\lambda_{\mathrm{R}}^{2}r}+\Big{(}\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{+},\lambda_{\mathrm{R}})\Big{)}-\Big{(}\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{-},\lambda_{\mathrm{R}})\Big{)}\;,\end{split}$
@@
Let $(H,\mu)$ be an insulator with $H\in\mathrm{AII}^{\hexagon}$. Then, for any $i,j=1,2$ the spin conductivity $\sigma^{\mathrm{s}}_{ij}=\sigma^{\mathrm{s}}_{ij}(H,\mu)$ can be written as
@@
$\widehat{G}(\mu;\mathbf{k})$ being the Fourier transform of the imaginary-time-ordered Green function:
$\widehat{G}(\mu;\mathbf{k}):=\big{(}H(k)-(\mu-\mathrm{i}k_{0})\mathbb{1}\big{)}^{-1}\;,\qquad\mu\in\mathbb{R}\;.$
@@
Note that $A_{\nu}$ is the connection associated with $\widehat{G}$. If there were $A_{\nu}$ in place of $A^{\mathrm{s}}_{\nu}$, the quantity on the r.h.s. of (4.23) would be proportional to the Pontryagin index associated with the connection $A_{\nu}$ [Col, Jack]. In general, the presence of the spin operator in $A^{\mathrm{s}}_{\nu}$ strips formula (4.23) of any geometrical interpretation, to the best of our knowledge.
@@
[…]
@@
By inspection, $H_{\mathrm{KM}}(k)$ and $S(m,\mathbf{q})$ are rational functions in $m$; therefore, $\widehat{G}(\mu_{c};\mathbf{k})$ and the terms $O_{m,\mathbf{q}}(1)$ are also rational functions in $m$. Since the latter terms are, in particular, bounded, they are also continuous in $m$. Since the singular part of $A_{\nu}$ and $A^{s}_{\nu}$ diverges as $(|\mathbf{q}|+|m|)^{-1}$, all the sub-leading contributions to $\operatorname{Tr}_{\mathbb{C}^{4}}\Big{(}A^{\mathrm{s}}_{i}(\mu_{c};\mathbf{k})A_{0}(\mu_{c};\mathbf{k})A_{j}(\mu_{c};\mathbf{k})\Big{)}$, namely those with at least one $O_{\mathbf{q},m}(1)$, are uniformly integrable over $\mathbb{T}^{2}_{+}$ as $m\to 0$. Therefore, by the dominated convergence theorem, these terms provide the same finite contribution to the limits $m\to 0^{\pm}$; in particular, they do not contribute to $\delta\sigma_{12}^{\mathrm{s}}$. For the same reason, we can extend the integration from $\mathbb{T}^{2}_{+}$ to the whole plane $\mathbb{R}^{2}$ of the singular contribution. Accordingly, we can write
@@
where
$\begin{split}\widetilde{\sigma}^{\mathrm{s}}_{12}(m)=-2\int_{\mathbb{R}^{3}}\frac{\mathrm{d}\mathbf{q}}{(2\pi)^{3}}\operatorname{Tr}_{\mathbb{C}^{4}}\Big{(}\widetilde{A}^{\mathrm{s}}_{1}(m;\mathbf{q})\widetilde{A}_{0}(m;\mathbf{q})\widetilde{A}_{2}(m;\mathbf{q})\Big{)}\;.\end{split}$ (4.26)
We now show that both limits in (4.25) exist (thus proving that the limits $\lim_{m\to 0^{\pm}}\sigma^{\mathrm{s}}_{12}(t,\lambda_{\mathrm{SO}},m,\lambda_{\mathrm{R}},r)$ exist) and compute their exact values. With the aid of a computer software [Wolf], we compute the trace in (4.26) and obtain
@@
The second term on the r.h.s. is odd in, e.g., $q_{1}$, so that it does not contribute to (4.26). Therefore,
$\begin{split}\delta\sigma_{12}^{\mathrm{s}}&=\Big{(}\lim_{m\to 0^{+}}-\lim_{m\to 0^{-}}\Big{)}\frac{3mt\widetilde{\alpha}\lambda_{\mathrm{SO}}^{3}}{4\pi^{3}}\int_{\mathbb{R}^{2}}\mathrm{d}q\int_{\mathbb{R}}\mathrm{d}q_{0}\big{(}q_{0}^{2}\alpha_{+}-2\mathrm{i}q_{0}m\lambda_{\mathrm{R}}^{2}+m^{2}\alpha_{-}+|q|^{2}\widetilde{\alpha}^{2}\big{)}^{-2}\;.\end{split}$
We compute the integral in $\mathrm{d}q_{0}$ by using that, for any $a,b,c\in\mathbb{R}$ such that $04ac+b^{2}>$ 0, we have $\int_{\mathbb{R}}\mathrm{d}z\big{(}az^{2}+\mathrm{i}bz+c\big{)}^{-2}=-4\pi a(4ac+b^{2})^{-\frac{3}{2}}$. Then, we note that for $a^{\prime},b^{\prime}>0$ we have $\int_{\mathbb{R}^{2}}\mathrm{d}x(a^{\prime}|x|^{2}+b^{\prime})^{-\frac{3}{2}}=2\pi\big{(}a^{\prime}b^{\prime 1/2}\big{)}^{-1}$, hence we have
$\begin{split}&\int_{\mathbb{R}^{2}}\mathrm{d}q\int_{\mathbb{R}}\mathrm{d}q_{0}\big{(}q_{0}^{2}\alpha_{+}-2\mathrm{i}q_{0}m\lambda_{\mathrm{R}}^{2}+m^{2}\alpha_{-}+|q|^{2}\widetilde{\alpha}^{2}\big{)}^{-2}\\ &\qquad=-4\pi\alpha_{+}\int_{\mathbb{R}^{2}}\mathrm{d}q\big{(}4|q|^{2}\alpha_{+}\widetilde{\alpha}^{2}+(2m\lambda_{\mathrm{SO}}^{2})^{2}\big{)}^{-\frac{3}{2}}=-\frac{\pi^{2}}{4\lambda_{\mathrm{SO}}^{2}\widetilde{\alpha}^{2}|m|}\;,\end{split}$

### v2

@@ 4.2. Phase diagram of the extended Kane–Mele model
First of all, we provide the dimerisation $H_{\mathrm{KM}}]_{{\mathcal{D}}_{i}}$. With abuse of notation, we let $T_{v}$ denote the translation operator on $\ell^{2}(\Gamma)$, whose definition is as in (2.7) with $\Gamma$ in place of $\mathcal{C}$. By straightforward computations, we obtain:
@@
$\sigma=(\sigma_{x},\sigma_{y},\sigma_{z})$ being the vector of Pauli matrices, and where $\{d_{j}\}_{j}$ and $\{a_{j}\}_{j}$ are the vectors introduced in (2.1) and (2.2). Let $H_{\mathrm{\sharp}}(k)$ denote the fibration of $H_{\mathrm{\sharp}}$ via the Bloch–Floquet transform at the momentum $k\in\mathbb{T}^{2}_{*}$, with $\sharp\in\{\text{NN},\text{SO},\text{W},\text{R1},\text{R2}\}$, having fixed the dimerisation in with $i=3$. Because $\big(\mathcal{F}\,T_{v}\psi\big)(k)=\mathrm{e}^{-\mathrm{i}k\cdot v}\,\mathcal{F}\psi(k)$ for any $v\in\Gamma$, it follows that
@@
and when $w=-\sigma\varepsilon\lambda_{\mathrm{SO}}$. Note that on the Brillouin torus $\mathbb{T}^{2}_{*}$, we have $k_{F}^{+}=k_{F}^{-}$.
@@
$k_{F}^{\pm}\cdot(d_{j}-d_{3})=\pm\frac{2\pi}{3}j\,(\mathrm{mod}\;2\pi)\;,\qquad k_{F}^{\pm}\cdot a_{j}=\mp\frac{2\pi}{3}\,(\mathrm{mod}\;2\pi)\;,$ (4.11)
@@
with $\cdot\,]_{+}:=\max\{\,\cdot\,,0\}$. Finally, defining the critical energy
@@
Note that the spin is not conserved and thus is not a good label for the energy bands. To compute the local internal gap $\Delta(k_{F}^{+})$, it is actually easier to first determine which bands are the external ones. Without loss of generality, we assume $\lambda_{\mathrm{SO}}>0$: otherwise, we could flip the sign of $w$ and $\lambda_{\mathrm{R}}$ to flip the sign of $H_{\mathrm{KM}}(k)$, the latter operation leaving the set of internal/external bands invariant. Note that $\mathcal{E}_{1,+}(k_{F}^{+})$ is always the positive external band. When $\mathcal{E}_{1,-}(k_{F}^{+})\geq\mathcal{E}_{2,-}(k_{F}^{+})$, the negative external band is $\mathcal{E}_{2,-}(k_{F}^{+})$, so that the local internal spectral gap is
@@
The condition $\mathcal{E}_{1,-}(k_{F}^{+})\geq\mathcal{E}_{2,-}(k_{F}^{+})$ is equivalent to $|\lambda_{\mathrm{R}}|\geq 2\sqrt{\lambda_{\mathrm{SO}}^{2}+|w|\lambda_{\mathrm{SO}}}$. In contrast, if $\mathcal{E}_{1,-}(k_{F}^{+})<\mathcal{E}_{2,-}(k_{F}^{+})$, $\mathcal{E}_{2,\pm}(k_{F}^{+})$ become the internal bands, and therefore $\Delta(k_{F}^{+})=2|w|$. The expression for $w_{c}$ and $\mu_{c}$ is a straightforward computation from $\Delta(k_{F}^{+})$, noticing that $\mathcal{E}_{2,+}(k_{F}^{+})$ is always an internal band. ∎
@@
We have thus shown that for $|\lambda_{\mathrm{R}}|$ small, the plot in Fig. 3 describes the quantum phase diagram of the model, consisting of three insulating phases separated by a semi-metallic one along the critical curves $w_{c}^{\pm}$. This phase diagram was already obtained for the Kane–Mele model ($r=0$) in the seminal work [KM1, Fig. 1, inset]. In particular, the central phase is a topological insulator with spin Chern number equal to $1$, whereas the outer regions have spin Chern number equal to 0. This can be directly inferred by continuity of the spin Chern number for insulators with respect to $\lambda_{\mathrm{R}}$, and thus by considering the case $\lambda_{\mathrm{R}}=0$, see (4.8), for which the spin Chern number is equal to the Chern number of $H_{+}$ in (4.8), which was computed in [Hal, Fig. 2 at $\phi=\pi/2$].
4.3. Lack of quantisation
Unlike the case of charge transport, the spin conductivity as introduced in Definition 3.12 is not universally quantised for time-reversal symmetric insulators.
@@ 4.3. Lack of quantisation
There exist insulators $(H,\mu)$ with $H$ belonging to $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$, time-reversal symmetric and almost conserving the spin (in the sense of Definition 3.20) such that $\sigma^{\mathrm{s}}_{12}(H,\mu)\notin\frac{1}{2\pi}\mathbb{Z}$. In particular, for such insulators it holds true that
[…]
@@
provided that the limit exists. Note that by Proposition 3.15 and Lemma 4.3 , it suffices to consider only $\sigma^{s}_{12}$, as $\sigma^{s}$ is an antisymmetric tensor. Note that we could equivalently study the jump across the lower critical line $w_{c}^{-}$, compare with (4.14).
@@
Assume the contrary, that is, $\sigma^{\mathrm{s}}_{12}(H,\mu)\in\frac{1}{2\pi}\mathbb{Z}$ for any insulator $(H,\mu)$ with $H$ time-reversal symmetric and belonging to $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. Then, $\delta\sigma_{12}^{\mathrm{s}}(t,\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}},r)\in\frac{1}{2\pi}\mathbb{Z}$ for any value of $t,\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}},r$, which is in contradiction with (4.22).
We are left with proving the lower bound in (4.18), the upper bound being stated in Theorem 3.22. Once more, we do this by considering the pair $(H_{\mathrm{KM}},\mu_{c})$ at fixed $t,\lambda_{\mathrm{SO}},r$ non-zero, at $\lambda_{\mathrm{R}}>0$ sufficiently small and at $w$ sufficiently close to $w_{c}^{+}(\lambda_{\mathrm{SO}},\lambda_{\mathrm{R}})$, that is, for $\neq 0, mm$ sufficiently small. Since all parameters but $m$ and $\lambda_{\mathrm{R}}$ are fixed, let us abridge (4.20) to $\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})$, and likewise $P^{\mathrm{sc}}=P^{\mathrm{sc}}(m)$, $w^{+}_{c}=w^{+}_{c}(\lambda_{\mathrm{R}})$, and $\delta\sigma_{12}^{s}=\delta\sigma_{12}^{s}(\lambda_{\mathrm{R}})$ (notice that $H^{\mathrm{sc}}=H_{\mathrm{KM}}|_{\lambda_{\mathrm{R}}=0}$).
@@
[…]
We now assume the negation of the claim and show that this leads to a contradiction. We first of all note that the negation of the claim implies that
$f(m,\lambda_{\mathrm{R}})=o(\lambda_{\mathrm{R}}^{2}).$
In fact, the negation of the claim is that for all $m\neq 0$, constants $C>0$, and open intervals $J\ni 0$, there exists at least one point $\lambda_{0}\in J\setminus\{0\}$, such that $|f(m,\lambda_{0})|<C\lambda_{0}^{2}$. Since this holds for any open interval $J\ni 0$, we can iterate the reasoning and find a sequence $(\lambda_{n})_{n}\in\mathbb{N}$ such that $\lambda_{n}\to 0$ and $|f(m,\lambda_{n})|<C\lambda_{n}^{2}$. By Remark 4.17, the limit $\lim_{\lambda_{\mathrm{R}}\to 0}f(m,\lambda_{\mathrm{R}})$ exists for any $m\neq 0$. As the constant $C>0$ can be chosen arbitrarily small,
$\lim_{\lambda_{\mathrm{R}}\to 0}f(m,\lambda_{\mathrm{R}})/\lambda_{\mathrm{R}}^{2}=0,$
that is, $f(m,\lambda_{\mathrm{R}})=o(\lambda_{\mathrm{R}}^{2})$.
It remains to show that the condition $f(m,\lambda_{\mathrm{R}})=o(\lambda_{\mathrm{R}}^{2})$ leads to a contradiction. To this end, we define
@@
$\mbox{$S$-Chern}(P^{\mathrm{sc}}(m))_{12}=\begin{cases}0&\qquad m>0\\ 1&\qquad m<0\;.\end{cases}$ (4.23)
@@
$\begin{split}g(m,\lambda_{\mathrm{R}})&=\frac{1}{2\pi}+\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})\\ &=\frac{1}{2\pi}+\delta\sigma_{12}^{s}(\lambda_{\mathrm{R}})+\Big(\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{+},\lambda_{\mathrm{R}})\Big)-\Big(\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{-},\lambda_{\mathrm{R}})\Big)\\ &=-\frac{1}{2\pi}\frac{\lambda_{\mathrm{R}}^{2}r}{2t\lambda_{\mathrm{SO}}-\lambda_{\mathrm{R}}^{2}r}+\Big(\sigma_{12}^{s}(m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{+},\lambda_{\mathrm{R}})\Big)-\Big(\sigma_{12}^{s}(-m,\lambda_{\mathrm{R}})-\sigma_{12}^{s}(0^{-},\lambda_{\mathrm{R}})\Big)\;.\end{split}$
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. Then, for any $i,j=1,2$ the spin conductivity $\sigma^{\mathrm{s}}_{ij}=\sigma^{\mathrm{s}}_{ij}(H,\mu)$ can be written as
@@
with $\widehat{G}(\mu;\mathbf{k})$ defined by
$\widehat{G}(\mu;\mathbf{k}):=\big(H(k)-(\mu-\mathrm{i}k_{0})\mathbb{1}\big)^{-1}\;.$
@@
Note that $A_{\nu}$ is the connection associated with $\widehat{G}$. If there were $A_{\nu}$ in place of $A^{\mathrm{s}}_{\nu}$, the quantity on the r.h.s. of (4.24) would be proportional to the Pontryagin index associated with the connection $A_{\nu}$ [Col, Jack]. In general, the presence of the spin operator in $A^{\mathrm{s}}_{\nu}$ strips formula (4.24) of any geometrical interpretation, to the best of our knowledge. Finally, note that $\widehat{G}(\mu;\mathbf{k})$ is the Fourier transform of the so-called imaginary-time-ordered Green function, see Appendix A.
@@
Note that, if the integrand in (4.25) were well-behaved, by the dominated convergence theorem we could take the limit inside the integral, thus proving that the limits $m\to 0^{\pm}$ exist and that $\delta\sigma_{12}^{\mathrm{s}}=0$. However, the integrand has an non-integrable divergence, which makes the existence of the limit non-trivial and $\delta\sigma_{12}^{\mathrm{s}}\neq 0$. To capture the singularity of the integrand, we introduce additional notation. By Proposition 4.9 we know that $(H_{\mathrm{KM}}(k)-\mu)^{-1}$ can be singular only at the Fermi points $k_{F}^{\pm}$, defined in (4.9); thus the only singular point in $\mathbb{R}\times\mathbb{T}^{2}_{+}$ for $\widehat{G}(\mu;\mathbf{k})$ is $\mathbf{k}_{F}^{+}:=(0,k_{F}^{+})\in\mathbb{R}\times\mathbb{T}^{2}_{+}$. We extract the most singular contribution from the Green function around $\mathbf{k}^{+}_{F}$. We use the notation $\mathbf{q}=(q_{0},q)\in\mathbb{R}\times\mathbb{R}^{2}$, and denote $|\,\cdot\,|$ the Euclidean norm. Recalling that $m=w-w_{c}^{+}$, see (4.19), and using the notation $O_{m,\mathbf{q}}(1)$ to denote terms that remain bounded as $|m|,|\mathbf{q}|\to 0$, we can write
@@
By inspection, $H_{\mathrm{KM}}(k)$ and $S(m,\mathbf{q})$ are rational functions in $m$; therefore, $\widehat{G}(\mu_{c};\mathbf{k})$ and the terms $O_{m,\mathbf{q}}(1)$ are also rational functions in $m$. Since the latter terms are, in particular, bounded, they are also continuous in $m$. Since the singular part of $A_{\nu}$ and $A^{s}_{\nu}$ diverges as $(|\mathbf{q}|+|m|)^{-1}$, all the sub-leading contributions to $\operatorname{Tr}_{\mathbb{C}^{4}}\Big(A^{\mathrm{s}}_{i}(\mu_{c};\mathbf{k})A_{0}(\mu_{c};\mathbf{k})A_{j}(\mu_{c};\mathbf{k})\Big)$, namely those with at least one $O_{\mathbf{q},m}(1)$, are uniformly integrable over $\mathbb{T}^{2}_{+}$ as $m\to 0$. Therefore, by the dominated convergence theorem, these terms provide the same finite contribution to the limits $m\to 0^{\pm}$; in particular, they do not contribute to $\delta\sigma_{12}^{\mathrm{s}}$. Accordingly, we can write
@@
where, denoting $\tilde{\mathbb{T}}_{+}^{2}:=\{q\in\mathbb{R}^{2}\,|\,k_{F}^{+}+q\in\mathbb{T}_{+}^{2}\}$,
$\begin{split}\widetilde{\sigma}^{\mathrm{s}}_{12}(m)=-2\int_{\mathbb{R}\times\tilde{\mathbb{T}}_{+}^{2}}\frac{\mathrm{d}\mathbf{q}}{(2\pi)^{3}}\operatorname{Tr}_{\mathbb{C}^{4}}\Big(\widetilde{A}^{\mathrm{s}}_{1}(m;\mathbf{q})\widetilde{A}_{0}(m;\mathbf{q})\widetilde{A}_{2}(m;\mathbf{q})\Big)\;.\end{split}$ (4.27)
We now show that both limits in (4.26) exist (and, as a matter of fact, also proving that the limits $\lim_{m\to 0^{\pm}}\sigma^{\mathrm{s}}_{12}(t,\lambda_{\mathrm{SO}},m,\lambda_{\mathrm{R}},r)$ exist) and compute their exact values. With the aid of a computer software [Wolf], we compute the trace in (4.27) and obtain
@@
The second term does not contribute to (4.27), because it is odd in $q_{1}$, whereas the integration domain $\tilde{\mathbb{T}}_{+}^{2}$ is even. Therefore,
$\begin{split}\delta\sigma_{12}^{\mathrm{s}}&=\Big(\lim_{m\to 0^{+}}-\lim_{m\to 0^{-}}\Big)\frac{3mt\widetilde{\alpha}\alpha_{-}^{2}}{16\lambda_{\mathrm{SO}}\pi^{3}}\int_{\mathbb{R}\times\tilde{\mathbb{T}}_{+}^{2}}\frac{\mathrm{d}\mathbf{q}}{\chi(m,\mathbf{q})^{2}}\\ &=\Big(\lim_{m\to 0^{+}}-\lim_{m\to 0^{-}}\Big)\frac{3mt\widetilde{\alpha}\lambda_{\mathrm{SO}}^{3}}{4\pi^{3}}\int_{\mathbb{R}^{3}}\mathrm{d}\mathbf{q}\big(q_{0}^{2}\alpha_{+}-2\mathrm{i}q_{0}m\lambda_{\mathrm{R}}^{2}+m^{2}\alpha_{-}+|q|^{2}\widetilde{\alpha}^{2}\big)^{-2}\;.\end{split}$
where in the second line we performed some algebraic manipulations and extended the integral to the whole $\mathbb{R}^{2}$ by dominated convergence, using that the integrand is uniformly integrable on $\mathbb{R}^{2}\setminus\tilde{\mathbb{T}}_{+}^{2}$ and continuous in $m$.
We compute the integral in $\mathrm{d}q_{0}$ using that, for any $a,b,c\in\mathbb{R}$ such that $4ac+b^{2}>0$, we have $\int_{\mathbb{R}}\mathrm{d}z\big(az^{2}+\mathrm{i}bz+c\big)^{-2}=-4\pi a(4ac+b^{2})^{-\frac{3}{2}}$. Then, we note that for $a^{\prime},b^{\prime}>0$ we have $\int_{\mathbb{R}^{2}}\mathrm{d}x(a^{\prime}|x|^{2}+b^{\prime})^{-\frac{3}{2}}=2\pi\big(a^{\prime}b^{\prime 1/2}\big)^{-1}$, hence we have
$\begin{split}&\int_{\mathbb{R}^{2}}\mathrm{d}q\int_{\mathbb{R}}\mathrm{d}q_{0}\big(q_{0}^{2}\alpha_{+}-2\mathrm{i}q_{0}m\lambda_{\mathrm{R}}^{2}+m^{2}\alpha_{-}+|q|^{2}\widetilde{\alpha}^{2}\big)^{-2}\\ &\qquad=-4\pi\alpha_{+}\int_{\mathbb{R}^{2}}\mathrm{d}q\big(4|q|^{2}\alpha_{+}\widetilde{\alpha}^{2}+(2m\lambda_{\mathrm{SO}}^{2})^{2}\big)^{-\frac{3}{2}}=-\frac{\pi^{2}}{\lambda_{\mathrm{SO}}^{2}\widetilde{\alpha}^{2}|m|}\;,\end{split}$